Odor Investigation

At 1249hrs Stations 1 & 2 was dispatched to Rotary Ln to assist Chief 720 for an odor investigation. Crews arrived and searched with the 4 gas and found only O2 @ 20.8 %. Chief 720 had the engine use the bumper line to flush out the storm sewer of any possible gases. Chief 720 had command and units were on scene for 15 mins. Engine 732 responded with (5)

Comments are closed.